OpenStudy (anonymous):

Solve the system (using substitution) x - 2y = -7 and 3x + 4y = -6. The y-value of the solution is:

OpenStudy (amistre64):

when solving by substitution we want to get a value for one variable in "terms" of the other. x-2y = -7 x = -7 +2y well use this for our value of "x" in the other equation. 3x +4y = -6 3(-7 +2y) + 4y = -6 -21 +6y +4y = -6 10y = -6+21 10y = 15 y = 15/10 = 1.5 the value of y should be 1.5 if I kept my wits about me :) better double check myself tho..

OpenStudy (amistre64):

x = -7 +2(1.5) x = -7 + 3 x=-4 -4 - 2(1.5) = -7 -4 -3 = -7 ..good there 3x + 4y = -6 3(-4) + 4(1.5) = -6 -12 +6 = -6 -6 = -6 ....its all good :)

OpenStudy (amistre64):

Does that make sense to you?

OpenStudy (anonymous):

Yes. That helped a lot. Thank you